OpenStudy (anonymous):

What is the exact value for the expression the square root of 171. − the square root of 19. + the square root of 76.? Simplify if possible.

OpenStudy (anonymous):

hint:171=9*19 76=4*19

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

171=9times19

OpenStudy (anonymous):

ok